# Accepted with Minor Revisions: Value of AI-Assisted Scientific Writing

### v1

@@ 3.4.2. Incentives
Reviewers earn a fixed participation fee of \$25 for participating in this study and completing all the tasks, irrespective of their decisions in the tasks. Leaning back onto the methodology of incentivized experiments, to increase data quality, we provide an additional bonus of \$8 to the reviewers whose evaluations align closely with those of their peers. One of the twenty abstracts is randomly selected, and two independent reviewers, alongside the reviewer-in-question, assess the same abstract pair using the slider scale described in Task 2. We then compute the median of the three ratings, and if a reviewer’s score falls within 10 points above or below this median, they receive the \$8 bonus. This peer-alignment incentive helps ensure careful, calibrated judgments and reduces random or inattentive responding. Additionally, they can earn a bonus of \$2 from Task 1. The first quartile of time taken for the reviewing task is also 1 hour and 8 minutes. However, the reviewers were not required to complete the task during one continuous sitting, and therefore, many took more than a day. Payments to reviewers were made through registered Amazon gift cards.
@@ 3.4.4. Reviewer Assignment
We adopted an automated pipeline to assign abstracts to recruited reviewers by mimicking the in-practice reviewer assignment mechanism in OpenReview, adopted by several CS conferences (stelmakh2023gold). In addition to topical fit, we also try to ensure that no reviewer can review more than one version of an edited abstract originating from the same original abstract to avoid contamination. First, we collect each reviewer’s scholarly profile using the Semantic Scholar APIs. Then we compute the semantic similarity between reviewer publications and base original abstracts using GPT-4o. We account for the reviewer’s citation count to weigh the similarity between an individual paper and submitted abstracts, so that we assign related abstracts closer to the reviewer’s most cited works. Finally, we solve the reviewer assignment as a constrained optimization problem by a minimum-cost flow, with constraints that each submitted abstract gets exactly three reviewers and each reviewer preferably does not review edit versions of the same original abstracts.
@@ 4.1.2. Heterogeneity through education
Figure 5. Patterns of editing effort in Study 1 (N = 495 abstracts). Left: Distribution of Levenshtein edit distance in two conditions: Human-noInfo (blue) and AI-noInfo (red). Black squares mark the sample means ± 95 % CI; the Welch unequal-variance test indicates a statistically reliable reduction in edits for AI abstracts (p = 0.0293). Middle: Mean edit distance broken down by the authors’ highest education level. Authors with undergraduate or graduate degrees make markedly larger edits to human text than to AI text, whereas doctorate-level editors show the opposite. Right: Scatter-plot of abstract’s perceived readability (0–100) vs the author’s confidence in their edits (0–100). The positive slopes show that authors feel more confident when an abstract reads more smoothly; readability, not AI, as the source explains authors’ higher confidence in the AI-noInfo condition.
@@ 5. Study B: Source Information Disclosure
The source of text, in recent times, could be human-written or AI-generated. While detection is often challenging, disclosing the source of a scientific text could be fundamental in shaping individual evaluation and interaction patterns with the text. Individuals have often reported reading AI-generated text more critically, anticipating errors, and focusing on the factual accuracy to encounter model hallucinations. In contrast, knowing that a text is human-produced can either invoke trust, or a reluctance to make additional edits. Individuals often feel a stronger sense of oversight when working with AI, or greater deference when editing human work. (buccinca2021trust; li2024does) provides evidence that transparent source information is crucial to altering calibration, confidence, and task performance.
@@ 5.1.4. Reviewer decisions
Figure 7. Reviewer decisions differ in AI conditions
We next examine whether the acceptance rate by the reviewers differs across treatment conditions. We model the final decision as a binary variable (1=1=Accept, 0=0=Reject) and compare acceptance rates across the three treatment contrasts, using both $2\times 2$ proportion tests and cluster-robust linear probability models with standard errors clustered by the original abstract. The acceptance rates remain nearly identical, with no statistical significance, across edited abstracts from human treatments: 0.730 in Human-noInfo versus 0.736 in Human-withInfo ($\Delta{=}{-}.006$). For edited abstracts from AI treatments, acceptance rates are 0.706 in AI-noInfo versus 0.773 in AI-withInfo, indicating a modest increase under disclosure ($\Delta{=}{-}.068$, p=0.081p{=}0.081). This suggests that edited versions of AI-generated abstracts, when authors know about the source, make definitive editing decisions that lead to higher acceptance, as shown in Figure 7. Keeping disclosure effect constant, comparing acceptance of edited abstracts from Human-withInfo (0.736) versus AI-withInfo (0.773) yields no significant difference. We further investigate how AI disclosure to authors may increase acceptance. Focusing on the marginal contrast in acceptance in edited abstracts from (AI-noInfo vs. AI-withInfo), we examine potential mediators reflecting editing effort, confidence, perceived readability, and AI beliefs. For each candidate mediator $M$, we estimate (i) $M{\sim}\text{Treatment}$ and (ii) $\text{FinalDecision}{\sim}\text{Treatment}+M$ using cluster-robust SEs by baseline abstract; the indirect effect was tested via Sobel’s $z$. Among all mediators, editing magnitude through edit distance exhibits the strongest pattern. Disclosure of AI authorship slightly shifts the volume of edits (p<.10p{<}.10), and higher editing effort predicts a higher likelihood of acceptance (p<.05p{<}.05), producing a marginally significant indirect effect. Authors make slightly different editing decisions when aware of AI authorship, and these editorial behaviors modestly enhance the final review outcome.
@@ 6. Study C: Quantitative and Qualitative Analyses of Edits
Stylistic guidance for scientific writing is grounded in empirical work on how readers cognitively process text. Psycholinguistics research shows that comprehension improves when information arrives in a predictable fashion. Readability of scientific writing improves when the topic position for context is known to the readers or the stress position for new or important information is expected (gopen1990science; halliday1967notes). Clarity in scientific prose improves when grammatical subjects mention real “characters” and verbs express actions, rather than nominalizing them (swales2014genre; williams2010style). We perform post-hoc quantitative and qualitative analyses of the edits to elicit editing strategies emerging from different treatment conditions.
@@ 6.2. Post-experiment Author Interviews
To align our quantitative understanding of how authors edit several treatment conditions, we conducted a series of semi-structured interviews following the main experiment. Studies 1 and 2 quantify how the implicit source and its disclosure moderate editing behavior, whereas these interviews attempt to elicit the reasoning behind such editing behavior. We interviewed 6 authors, randomly sampled across all four treatment conditions, who successfully completed the editing task. Each interview, conducted over Zoom, lasted for 15 minutes, for which they were compensated with a \$10 flat fee. Post-hoc thematic analysis of the transcripts reveals how authors adopted several editing strategies and their perception of AI-assistance in scientific writing.
@@
7. Discussions and Conclusion
We offer a systematic evaluation of AI-assisted scientific writing through a randomized controlled trial simulating a peer review process. Our results show that authorial behavior is significantly influenced by both the implicit source of an abstract and the disclosure of it: authors made the most revisions to human-written texts when source information was withheld, whereas AI-generated abstracts received mildly more edits when attribution was not disclosed. However, reviewer decisions were not significantly affected by the abstract’s source. Instead, acceptance outcomes were primarily driven by the degree of meaningful edits. Our analysis suggests that careful edits on the AI-generated abstract can yield higher, if not equal, acceptance as compared to abstracts that are human-authored.
@@ 7. Discussions and Conclusion
This work establishes a foundation for evaluating AI-generated scientific writing in realistic incentivized authorial and review settings. Future research should extend this methodology to other domains beyond computer science to understand domain-specific interactions with AI. Investigating collaborative workflows such as co-writing between multiple authors with AI support would uncover both potential and pitfalls in the use of AI assistance. Additionally, scaling to full-paper evaluations and incorporating multimodal inputs (e.g., tables, figures) would provide a more comprehensive assessment of LLM utility in scientific communication. Finally, data created using such studies could also be used to train models to conduct edits to further improve text quality in the domain of science.
@@
Looking beyond a standalone experiment, longitudinal studies could help us understand how repeated use of AI-assisted writing affects authors’ skill development, stylistic habits, and confidence over time, as well as how it reshapes norms around peer review and scientific communication. Finally, we raise an important point: ethical considerations around transparency, attribution, and responsible use of AI in research are understudied and can have welfare-altering impacts. Our study highlights the importance of academic experiences for authors, which play an important role in understanding and effectively utilizing the benefits of AI assistance. At a large scale, we must expand training programs, which are lacking, to inform novice researchers about the socially beneficial use of AI in science. Community-driven investigations of responsible use of AI in science thus must uncover downstream effects on scientific rigor, reproducibility, and community norms to optimally reap the democratic benefits of AI.
@@
We used GPT-4o to generate AI-generated abstracts central to our experiments. Similarly, GPT-4o was used for assigning reviewers to relevant abstracts for review, again critical for our experiments.
Appendix A Editing Interface
Figure 9. The above image illustrates how the authors view the abstracts. The left non-editable, copy-paste disabled panel includes the excerpts from a selected research paper, and the right panel includes the embedded firepad author where the authors can make changes at the keystroke level.
Appendix B Screening Questions
Figure 10. Screening questions assess the authors’ understanding of task.
Appendix C Comments
@@
Appendix D Distinguishing AI and Human Texts
@@
Appendix E List of Papers
Table 3. List of papers we used and the venues where they were published.
@@ Appendix E List of Papers
Appendix F Prompts and Output
GPT-4o prompt for extracting research plan and findings from published papers Given this research paper, please extract content that will be useful for understanding the concept, background, findings, and result of the paper. For EVERY sentence extracted, please provide the page number from the information is taken in a square bracket as [page number]. DO NOT PRODUCE ANYTHING THAT IS NOT WRITTEN IN THE PAPER. DO NOT LEAVE OUT ANY RELEVANT INFORMATION. EXTRACT THE EXACT SENTENCE FROM THE PAPER. EXCLUDE ANY CONTENTS FROM THE ABSTRACT OR CONTENT BEFORE INTRODUCTION. DO NOT INCLUDE ANY EQUATIONS. THE LENGTH OF THE CONTENT SHOULD BE AROUND 1000 WORDS. DO NOT INCLUDE ANY REFERENCE TO FIGURES OR TABLES. Give exact sentences without any formatting or dividing sections. Use linebreaks. The extracted content should be sufficient to write the abstract.
Figure 13. GPT-4o Prompt we used to extract research findings from papers. The output of the model (after verification) is used as the excerpt to generate abstract.
An Example of extracted research plan and findings from published papers The capacity to generate and manipulate digital 3D objects lies at the core of a multitude of applications related to the entertainment and media industries, virtual and augmented reality and healthcare [1]. However, achieving fine control in mesh manipulation necessitates the learning of a disentangled representation of 3D shapes which is still an open research problem [1]. Recently proposed methods based on Graph Convolutional Network (GCN)-based Auto-Encoders (AEs) have demonstrated impressive performance in dimensionality reduction but typically learn a highly entangled latent space making them unsuitable for detailed shape manipulation [1]. Additionally, despite having a low parameter count, these methods struggle to handle high-resolution meshes, limiting their applicability [1]. Few works have dealt with the disentanglement of local identity attributes, however these methods still rely on GCNs and opt for controlling manipulations through the state of the latent code which is partitioned and assigned to predefined object regions [1]. Using the latent code to drive shape manipulation requires the use of explicit optimization objectives to learn a disentangled latent space [1]. Moreover, partitioning its state is critically suboptimal for learning compressed representations of 3D objects [1]. We propose a different paradigm which does not involve partitioning the latent code or relying on its state to drive changes in shape, resulting in state-of-the-art (SOTA) disentanglement and reconstruction capabilities in a unified architecture [1]. Instead, we use a global latent code for 3D object unconditional generation and utilise additional inputs to jointly train our generative model to locally overwrite the latent encoded geometry [1]. We present the Locally Adaptive Morphable Model (LAMM), a general framework for manipulating the geometry of registered meshes [1]. To the best of our knowledge, this is the first method that allows direct shape control with a single forward pass [1]. Applied on human 3D heads, LAMM exhibits SOTA disentanglement properties and allows for very fine geometric control over both facial identities and expressions [1]. Our models, trained for manipulation, concurrently exhibit SOTA performance in mesh dimensionality reduction compared against methods trained exclusively on this task [1]. As a result, a single model can be used to generate entirely new shapes and apply both localized and global modifications to their geometry [1]. We show how our framework can leverage direct control as a primitive to achieve higher level editing operations such as region swapping and sampling [1]. $\cdots$
@@ Appendix F Prompts and Output
GPT-4o prompt for generating abstracts from extracted research plan and findings You will be presented with an excerpt from research papers in Computer Science. I have curated these excerpts from research papers which have been published in top-tier peer-reviewed articles. You goal is to write an abstract. What is an abstract? An abstract is a short summary of completed research. It is intended to describe the work without going into great detail. Abstracts should be self-contained and concise, explaining your work as briefly and clearly as possible. An abstract should be able to stand independently from the research paper and still tell the reader something significant. The most important function of an abstract is to help a reader decide if he or she is interested in reading your entire publication. An effective abstract will contain several key features: Motivation or problem statement: Why is the research/argument important? What practical, scientific, theoretical or artistic gap is the project filling? Methods/procedure/approach: What did the researcher actually do to get your results? (e.g. analyzed 3 novels, completed a series of 5 oil paintings, interviewed 17 students) Results/findings/product: After completing the above procedure, what did the researcher learn/invent/create? Conclusion/implications: What are the larger implications of the findings, especially for the problem/gap identified previously? Why is this research valuable? Keep the abstract short: A general rule of abstract length is 150-200 words. Do not add any new information: If something doesn’t appear in the input excerpt, then don’t put it in the abstract. An abstract is supposed to convey scientific findings, so they have to be precise and factual. Please don’t embellish any results or findings. Excepts for writing the abstract: [Excerpt]
@@
An example of generated abstract from extracted research plan and findings The generation and manipulation of 3D meshes are pivotal for applications across entertainment, virtual and augmented reality, and healthcare. However, existing methods struggle to achieve fine-grained, spatially disentangled control of 3D shapes, often relying on Graph Convolutional Network (GCN)-based auto-encoders (AEs) that produce highly entangled latent spaces and face scalability issues with high-resolution meshes. We propose the Locally Adaptive Morphable Model (LAMM), a novel framework that circumvents these limitations by enabling direct, localized mesh manipulation without partitioning the latent code or introducing explicit disentanglement objectives. LAMM employs fully connected layers and a global receptive field to respect the distinct semantics of registered mesh vertices, allowing for both unconditional shape generation and fine-grained editing via direct control inputs. Extensive experiments on 3D human head and hand meshes demonstrate that LAMM achieves state-of-the-art performance in both disentanglement and reconstruction while scaling efficiently to high-resolution data. Our model surpasses existing methods in terms of memory efficiency and inference speed, offering practical benefits such as faster CPU-based execution and reduced resource requirements. LAMM’s capabilities democratize access to advanced 3D modeling techniques, enabling applications ranging from detailed facial identity editing to region swapping and geometric sampling.
@@
An example of human written abstract We present the Locally Adaptive Morphable Model (LAMM), a highly flexible Auto-Encoder (AE) framework for learning to generate and manipulate 3D meshes. We train our architecture following a simple self supervised training scheme in which input displacements over a set of sparse control vertices are used to overwrite the encoded geometry in order to transform one training sample into another. During inference, our model produces a dense output that adheres locally to the specified sparse geometry while maintaining the overall appearance of the encoded object. This approach results in state-of-the-art performance in both disentangling manipulated geometry and 3D mesh reconstruction. To the best of our knowledge LAMM is the first end-to-end framework that enables direct local control of 3D vertex geometry in a single forward pass. A very efficient computational graph allows our network to train with only a fraction of the memory required by previous methods and run faster during inference, generating 12k vertex meshes at ¿60fps on a single CPU thread. We further leverage local geometry control as a primitive for higher level editing operations and present a set of derivative capabilities such as swapping and sampling object parts. Code and pretrained models can be found at https://github.com/michaeltrs/LAMM.

### v2

@@ 3.4.2. Incentives
Reviewers earn a fixed participation fee of \$25 for participating in this study and completing all the tasks, irrespective of their decisions in the tasks. Leaning back onto the methodology of incentivized experiments, to increase data quality, we provide an additional bonus of \$8 to the reviewers whose evaluations align closely with those of their peers. One of the twenty abstracts is randomly selected, and two independent reviewers, alongside the reviewer-in-question, assess the same abstract pair using the slider scale described in Task 2. We then compute the median of the three ratings, and if a reviewer’s score falls within 10 points above or below this median, they receive the \$8 bonus. This peer-alignment incentive helps ensure careful, calibrated judgments and reduces random or inattentive responding. Unlike Task 1 incentive for authors, reviewers earn a bonus of \$2 from Task 1, task remaining identical. The first quartile of time taken for the reviewing task is also 1 hour and 8 minutes. However, the reviewers were not required to complete the task during one continuous sitting, and therefore, many took more than a day. Payments to reviewers were made through registered Amazon gift cards.
@@ 3.4.3. Procedures and Reviewer Recruitment
Figure 5. Patterns of editing effort in Study 1 (N = 495 abstracts). Left: Distribution of Levenshtein edit distance in two conditions: Human-noInfo (blue) and AI-noInfo (red). Black squares mark the sample means ± 95 % CI; the Welch unequal-variance test indicates a statistically reliable reduction in edits for AI abstracts (p = 0.0293). Middle: Mean edit distance by the authors’ highest education level. Authors with undergraduate or graduate degrees make markedly larger edits to human text than to AI text, whereas doctorate-level editors show the opposite. Right: Scatter-plot of abstract’s perceived readability (0–100) vs the author’s confidence in their edits (0–100). The positive slopes show that authors feel more confident when an abstract reads more smoothly; readability, not AI, as the source explains authors’ higher confidence in the AI-noInfo condition.
@@ 3.4.4. Reviewer Assignment
We adopted an automated pipeline to assign abstracts to recruited reviewers by mimicking the in-practice reviewer assignment mechanism in OpenReview, adopted by several CS conferences (Stelmakh et al., 2023). In addition to topical fit, we also try to ensure that no reviewer can review more than one version of an edited abstract originating from the same original abstract to avoid contamination. First, we collect each reviewer’s scholarly profile using the Semantic Scholar APIs. Then we compute the semantic similarity between reviewer publications and base original abstracts using GPT-4o. We account for the reviewer’s citation count to weigh the similarity between an individual paper and submitted abstracts, so that we assign related abstracts closer to the reviewer’s most cited works. Finally, we solve the reviewer assignment as a constrained optimization problem by a minimum-cost flow, with constraints that each submitted abstract gets exactly three reviewers and each reviewer preferably does not review edit versions of the same original abstracts.
@@ 5. Study B: Source Information Disclosure
The source of text, in recent times, could be human-written or AI-generated. While detection is often challenging, disclosing the source of a scientific text could be fundamental in shaping individual evaluation and interaction patterns with the text. Individuals have often reported reading AI-generated text more critically, anticipating errors, and focusing on the factual accuracy to encounter model hallucinations. In contrast, knowing that a text is human-produced can either invoke trust, or a reluctance to make additional edits. Individuals often feel a stronger sense of oversight when working with AI, or greater deference when editing human work. (Buçinca et al., 2021; Li et al., 2024) provides evidence that transparent source information is crucial to altering calibration, confidence, and task performance.
@@ 5.1.4. Reviewer decisions
We next examine whether the acceptance rate by the reviewers differs across treatment conditions. We model the final decision as a binary variable (1=1=Accept, 0=0=Reject) and compare acceptance rates across the three treatment contrasts, using both $2\times 2$ proportion tests and cluster-robust linear probability models with standard errors clustered by the original abstract. The acceptance rates remain nearly identical, with no statistical significance, across edited abstracts from human treatments: 0.730 in Human-noInfo versus 0.736 in Human-withInfo ($\Delta{=}{-}.006$). For edited abstracts from AI treatments, acceptance rates are 0.706 in AI-noInfo versus 0.773 in AI-withInfo, indicating a modest increase under disclosure ($\Delta{=}{-}.068$, p=0.081p{=}0.081). This suggests that edited versions of AI-generated abstracts, when authors know about the source, make definitive editing decisions that lead to higher acceptance, as shown in Figure 7.
Keeping disclosure effect constant, comparing acceptance of edited abstracts from Human-withInfo (0.736) versus AI-withInfo (0.773) yields no significant difference. We further investigate how AI disclosure to authors may increase acceptance. Focusing on the marginal contrast in acceptance in edited abstracts from (AI-noInfo vs. AI-withInfo), we examine potential mediators reflecting editing effort, confidence, perceived readability, and AI beliefs. For each candidate mediator $M$, we estimate (i) $M{\sim}\text{Treatment}$ and (ii) $\text{FinalDecision}{\sim}\text{Treatment}+M$ using cluster-robust SEs by baseline abstract; the indirect effect was tested via Sobel’s $z$. Among all mediators, editing magnitude through edit distance exhibits the strongest pattern. Disclosure of AI authorship slightly shifts the volume of edits (p<.10p{<}.10), and higher editing effort predicts a higher likelihood of acceptance (p<.05p{<}.05), producing a marginally significant indirect effect. Authors make slightly different editing decisions when aware of AI authorship, and these editorial behaviors modestly enhance the final review outcome.
@@ 6. Study C: Quantitative and Qualitative Analyses of Edits
Stylistic guidance for scientific writing is grounded in empirical work on how readers cognitively process text. Psycholinguistics research shows that comprehension improves when information arrives in a predictable fashion. Readability of scientific writing improves when the topic position for context is known to the readers or the stress position for new or important information is expected (Gopen and Swan, 1990; Halliday, 1967). Clarity in scientific prose improves when grammatical subjects mention real “characters” and verbs express actions, rather than nominalizing them (Swales, 2014; Williams and Bizup, 2010). We perform post-hoc quantitative and qualitative analyses of the edits to elicit editing strategies emerging from different treatment conditions.
@@ 6.1.1. Stylistic Shifts Across Treatments
6.2. Edits and AI-detection
To measure the significance of the edits and how AI being the source (perceived or disclosed) moderates editing behavior, we use Pangram [^18] https://www.pangram.com/, an automated AI detection tool, to find cases where author’s edit made the edited abstract more human-written. We can only analyze cases where the provided abstract was AI-generated (i.e., AI-noInfo and AI-withInfo) and use AI detection to label if the edited abstract still can be perceived as “AI-generated.”
We first examine treatment-level differences in the probability that an edited abstract is no longer classified as AI. Authors edited abstracts in the disclosure condition (AI-withInfo) to make the abstract more human-like (23% cases, $N=225$), i.e., can no longer be classified as AI-generated, as compared to no-disclosure condition (AI-noInfo; 18% cases, $N=225$), although this difference is not statistically significant. To characterize this behavior through character-level edits, we find that disclosure significantly increases editing intensity, making edited abstracts more human-like. Conclusively, character-level edit distance emerges as a strong predictor of AI detection outcomes ($p<0.001$).
@@ 6.3. Post-experiment Author Interviews
To align our quantitative understanding of how authors edit several treatment conditions, we conducted a series of semi-structured interviews following the main experiment. Studies 1 and 2 quantify how the implicit source and its disclosure moderate editing behavior, whereas these interviews attempt to elicit the reasoning behind such editing behavior. We interviewed 6 authors across all four treatment conditions who successfully completed the editing task. Each interview, conducted over Zoom, lasted for 15 minutes, for which they were compensated with a \$10 flat fee. Post-hoc thematic analysis of the transcripts reveals how authors adopted several editing strategies and their perception of AI-assistance in scientific writing.
@@
7. Discussion
Here we summarize our key findings connecting studies A (Section 4), B (Section 5), and C (Section 6), grounding in prior work and with possible design implications.
Perceived or disclosed source acts as a reputational signal.
Study A shows fewer edits to AI-generated abstracts when the source is undisclosed. Study B shows disclosure (weakly) reverses this pattern, especially reducing edits to human-written abstracts. This result is consistent with prior work showing that source labels activate algorithm aversion and source-identity bias even when content quality is held constant (Dzindolet et al., 2003). Authors explicitly reference deference, accountability, and authority in the interviews. This aligns with evidence in prior work that perceived authorship and authority trigger deference and reduced intervention in evaluative and collaborative settings (Zhu et al., 2025; Fiedler and Döpke, 2025). Collaborative systems that foster AI assistance in scientific writing then must reveal origin of text (e.g., through watermarking) that can assign the appropriate accountability of the authors.
Readability and stylistic edits on AI writing.
Authors express higher perceived readability for AI-generated abstracts in no-disclosure conditions. Still, they make meaningful stylistic edits such as reducing nominalization and improving cohesion. Interview evidence additionally confirms simplification strategies for AI text, also tying back to perceived or disclosed source, acting as a reputational signal. These readability and stylistic patterns reflect broader findings from recent corpus studies showing that AI-generated texts often exhibit distinct linguistic profiles such as more uniform sentence structure and enhanced surface-level coherence (Xu, 2025). AI writing assistants could improve these obvious sloppiness to meaningfully contribute in academic writing.
Expertise mediates trust on AI writing.
Authors with PhD-level experience exhibit completely opposite editing behavior, editing AI-generated abstracts heavily. Mixed-effects models show variance driven by author characteristics rather than abstract quality, which was corroborated by interview narratives reflecting a stronger critical posture among experienced authors. Prior studies similarly show that experts are less susceptible to fluency-driven trust in AI-generated text and are more likely to detect limitations or stylistic artifacts, whereas less experienced users rely more heavily on surface-level coherence as a quality signal (Coutrix and Prost, 2024; Chakrabarty et al., 2024a).
Asymmetric trust on human and AI
Authors report equal responsibility in checking correctness across AI and human text, but more weights put when the text is AI-generated, which was revealed consistently during interviews. Indeed, source disclosure influences how authors edit, hence their trust allocation. When the source is disclosed as AI, we find a strong heterogeneity: authors willing to use AI reduce edits to AI-generated abstracts. On the other hand, with the source being disclosed as human, it substantially reduces the overall edit for human-written abstracts. But, in general, even though editing behavior changes across source information conditions depending on the disclosure, it does not translate into systematic differences in reviewer acceptance. This divergence suggests that disclosure primarily affects authors’ perception of reputation or a manifestation of algorithm aversion, rather than the substantive quality of the final text.
Limitations and future work
@@ 7. Discussion
Like most experimental studies, our setup also has some limitations which could be addressed in a future work. Though elaborate, our setup is lab-based: a similar study could be done in a real conference setup, where we could observe author behavior with natural incentives. Similarly, capturing editing behavior with real-time AI help, contrary to our offline setup where AI abstracts are already provided. Additionally, confidence and readability plays an important role in describing author’s editing behavior, but we do not incentivize to provide true beliefs to keep the study simple. Finally, This work establishes a foundation for evaluating AI-generated scientific writing in realistic incentivized authorial and review settings. Future research should extend this methodology to other domains beyond computer science to understand domain-specific interactions with AI. Investigating collaborative workflows such as co-writing between multiple authors with AI support would uncover both potential and pitfalls in the use of AI assistance. Additionally, scaling to full-paper evaluations and incorporating multimodal inputs (e.g., tables, figures) would provide a more comprehensive assessment of LLM utility in scientific communication. Finally, data created using such studies could also be used to train models to conduct edits to further improve text quality in the domain of science.
8. Conclusion
We offer a systematic evaluation of AI-assisted scientific writing through a randomized controlled trial simulating a peer review process. Our results show that authorial behavior is significantly influenced by both the implicit source of an abstract and the disclosure of it: authors made the most revisions to human-written texts when source information was withheld, whereas AI-generated abstracts received more extensive edits when attribution was disclosed. However, reviewer decisions were not significantly affected by the abstract’s source. Instead, acceptance outcomes were primarily driven by the degree of meaningful edits. Our analysis suggests that careful edits on the AI-generated abstract can yield higher, if not equal, acceptance as compared to abstracts that are human-authored.
@@ 8. Conclusion
Looking beyond a standalone experiment, longitudinal studies could help us understand how repeated use of AI-assisted writing affects authors’ skill development, stylistic habits, and confidence over time, as well as how it reshapes norms around peer review and scientific communication. Finally, we raise an important point: ethical considerations around transparency, attribution, and responsible use of AI in research are understudied and can have welfare-altering impacts (Hazra et al., 2025). Our study highlights the importance of academic experiences for authors, which play an important role in understanding and effectively utilizing the benefits of AI assistance. At a large scale, we must expand training programs, which are lacking, to inform novice researchers about the socially beneficial use of AI in science. Community-driven investigations of responsible use of AI in science thus must uncover downstream effects on scientific rigor, reproducibility, and community norms to optimally reap the democratic benefits of AI.
@@
We used GPT-4o to generate AI-generated abstracts central to our experiments. Similarly, GPT-4o was used for assigning reviewers to relevant abstracts for review, again critical for our experiments. We use generative AI-based tools (e.g., Grammarly) to make grammatical checks and grammatical edits for our manuscript.
Acknowledgements.
This material is based upon work supported by the Ai2 Faculty Research Award and the OSU Kirwan Institute. We are grateful to Tuhin Chakraborty, Joseph Chang, Pao Siangliulue, and Peter Clark for feedback on early drafts of this paper. We also thank the anonymous reviewers for their insightful comments.
Appendix A Additional Tables and Figures
Figure 9. This is how the authors view the abstracts. The left non-editable, copy-paste disabled panel includes the excerpts from a selected research paper, and the right panel is the embedded Firepad editor where the authors can make changes.
Figure 10. Screening questions to assess the authors’ understanding of the task.
A.1. Author Comments
@@
A.2. Distinguishing AI and Human Texts
@@
Table 3. List of selected papers with venues & years
@@
Figure 13. GPT-4o Prompt we used to extract research findings from papers. The output of the model (after verification) is used as the excerpt to generate the abstract.